OpenStudy (anonymous):

evaluate the integral (32x^3dx)/sqrt(10-x^4), using u = 10 - x^4

OpenStudy (anonymous):

\[=32 \int\limits \frac{x^3}{\sqrt(10-x^4)} dx\] let : \[u = 10-x^4\] \[du = -4x^3 dx\] so : \[= 32(\frac{-1}{4}) \int\limits \frac{1}{\sqrt(u)} du \] \[= -8 \int\limits u^{\frac{-1}{2}} du\] \[= -8 [ 2\sqrt(u)] + c = -16 \sqrt(10-x^4) + c\] Hope it's clear now ^_^